A cylinder with a movable piston contains 2.00 g of helium, He, at room temperature. More helium was added to the cylinder and the volume was adjusted so that the gas pressure remained the same. How many grams of helium were added to the cylinder if the volume was changed from 2.00 L to 4.10 L ? (The temperature was held constant.)
1
Expert's answer
2019-11-07T08:04:12-0500

pV=nRT

p=n1RT/V = 0.5 x 8.314 x 298 / 2 = 619.393 kpa.

n2= pV/RT= 619.393 x 4 / 8.314 x 298 = 1 mol.

m=nxMr = 1 x 4 = 4 gr.